OpenStudy (anonymous):

Brenda is training to be a volunteer at her library. She wonders what the probability is that the first item she'll get to check out for a customer will be a video. So, on her first day of training, she tracks what items are checked out. The results are shown here in a tally chart. Books: 37 Videos: 24 Mgazines: 19 Use Brenda's experimental data to estimate the probability that her first item to check out will be a video. Show and explain your work.

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

anyone?

thomaster (thomaster):

The total is 37+24+19= 80 items 24 items are video's Do you know how to calculate the percentage chance?

OpenStudy (nincompoop):

we'll see if opcode helps you first

OpenStudy (anonymous):

no i was wondering that too

OpenStudy (nincompoop):

if we showed and explained to you how, that'll be our work and not yours :(

OpenStudy (anonymous):

would i have to divide?

thomaster (thomaster):

just divide the part (videos) by the total and multiply by 100% :)

OpenStudy (anonymous):

okay so you are telling me to divide 24 by 80 ?

thomaster (thomaster):

\(\large\dfrac{24}{80}*100\%\)

OpenStudy (anonymous):

24/80 =0.3 now what do i do?

thomaster (thomaster):

multiply by 100

OpenStudy (anonymous):

so then my answer is 30 ?

thomaster (thomaster):

Yes, the probability is 30% that she'll check out a video first

OpenStudy (anonymous):

ok thanks so much you guys! medal to the both of you!

OpenStudy (anonymous):

...